Peter said "I am three times as old as my brother, but in 6 years time I will only be twice as old as him". If Peter's brother's age now is x years, how old is Peter in terms of x ? How old will they be in 6 years time? write down an equation and solve it to find there ages now.
Answer by stanbon(75887)

Peter said "I am three times as old as my brother, but in 6 years time I will only be twice as old as him". If Peter's brother's age now is b years, how old is Peter in terms of b ? How old will they be in 6 years time?
------
Equations:
p = 3b
p+6 = 2(b+6)
----
Substitute for "p" and solve for "b":
3b+6 = 2b + 12
b = 6 years (Brother's age now)
-------
how old is Peter in terms of b ?
p = 3b = 18 years
--------------------------
How old will they be in 6 years time?
Peter will be 24 yrs
Brother will be 12 yrs.
============================
Cheers,
Stan H.
============================